Belinda had &20,000 to invest. She invested part of it at 10% and the remainder at 12%. If her income from the two investments was $2160, then how much did she invest at each rate?
Answer by richwmiller(17219)

x+y=20000,
0.1*x+0.12*y=2160
x=20000-y
0.1*(20000-y)+0.12*y=2160
2000-0.1y+0.12*y=2160
0.02*y=160
y= 8000
x=20000-y
x=12000
check
0.1*12000+0.12*8000=2160
1200+960=2160
2160=2160
ok